OpenStudy (anonymous):

Find the direct variation equation which passes through (0, 0) and (4, 1).

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (1-0)/(4-0) m = 1/4 -------------------------------- y = mx+b 0 = (1/4)(0)+b 0 = 0+b 0 = b b = 0 y = mx+b y = (1/4)x+0 y = (1/4)x Equation is \[\Large y = \frac{1}{4}x\]